A ball of mass 2 kg is projected up in the air from ground level with a speed 20 ms-1. it experiences constant air resistance R. It returns ground level with a speed of 15 ms-1 find R
1
Expert's answer
2020-02-17T09:08:15-0500
"a_1=\frac{R}{m}+g"

"v_1^2-0^2=2a_1h"

"a_2=g-\frac{R}{m}"

"v_2^2-0^2=2a_2h"

Thus,


"\frac{a_1}{a_2}=\frac{\frac{R}{m}+g}{g-\frac{R}{m}}=\frac{v_1^2}{v_2^2}"

"\frac{\frac{R}{2}+9.8}{9.8-\frac{R}{2}}=\frac{20^2}{15^2}"

"R=5.5\ N"